OpenStudy (anonymous):

A 400 ohms resistor is connected in series with a 2358pF capacitor across a 12V A.C. supply. determine the supply frequency, if the current flowing in the circuit is 24mA

OpenStudy (anonymous):

What is the difficulty you are having with this question ? Do you know the impedance of a capacitor ?

OpenStudy (radar):

To begin with, calculate the impedance using Z=E/I. After that is known you then use the formula for Z which is:\[Z ^{2}=R ^{2}+X ^{2}_{c}\]After getting the capacitance Xc solve for frequency using:\[x _{c}=1/(2pifC)\]Then solve for f.

OpenStudy (arnavguddu):

Calculate Z from Z = AC voltage / AC current. Now Z^2 = R^2 + Xc^2.... Substitute Z, R and so Xc^2 = Z^2 - R^2 Next Xc = 1/(2 pi f C). So f = 1/(2 pi Xc C). substitute values and done!!!